# PPO-ACT: Proximal Policy Optimization with Adversarial Curriculum Transfer for Spatial Public Goods Games

### v1

@@ 3.2 PPO-ACT with half-and-half initialization
Figure 2: Evolution of cooperative behavior in PPO-ACT framework. (a,c) Temporal evolution under enhancement factors $r=3.0$ (a) and $r=4.0$ (c) showing the complete Phase 1 + Phase 2 training process, with strategy space snapshots (white: cooperator, black: defector) at iterations $\{0,10,100,1000,10000\}$. (b,d) Corresponding Phase 2-only results for $r=3.0$ (b) and $r=4.0$ (d) with snapshots at $\{0,1,10,100,1000\}$ iterations. Initial conditions: defectors (red) occupy the upper grid half, cooperators (blue) lower half. All panels share identical color mapping and spatial scale bars.
@@ 3.3 PPO-ACT with bernoulli random initialization
(b) Phase 1
Figure 3: Evolution of cooperative behavior under Bernoulli initialization ($p=0.5$). (a) Complete training process with Phase 1 ($r=5.0$, 0-999 iterations) and Phase 2 ($r=4.0$, from iteration 1000). Snapshots captured at $t=[0,10,100,1000,10000]$ (white: cooperator, black: defector). (b) Phase 2-only dynamics after strategy reinitialization. Snapshots at $t=[0,1,10,100,1000]$ demonstrate immediate cooperation fixation.
@@ 3.4 PPO-ACT with all-defectors initialization
Figure 4: Initial agent strategies were set with all agents as defectors. (a) Full process: Phase 1 ($r=5.0$, 0-999 iterations) and Phase 2 ($r=4.8$, from iteration 1000). Snapshots at $t=[0,10,100,1000,10000]$. (b) Phase 2-only process. Snapshots at $t=[0,1,10,100,1000]$. White indicates cooperators, and black denotes defectors.
@@ 3.5 Comparative analysis of algorithms
Figure 5: Comparative analysis of PPO-ACT, PPO, Q-learning, and Fermi update rule: Temporal evolution curves of cooperators (blue) and defectors (red) at $r=4.0$, with corresponding snapshots of cooperators (white) and defectors (black). The initialization strategy places defectors in the upper half and cooperators in the lower half. Snapshots are shown chronologically from left to right (0, 10, 100, 1000, and 10000 iterations).
@@ 3.6 Comparative analysis of algorithms across different $r$
Figure 6: Performance comparison of PPO-ACT, PPO, Q-learning, and Fermi update rule across different r-values: cooperators (blue) and defectors (red). Initial agent strategies were set with all defectors in the upper grid half and all cooperators in the lower half.
@@
PPO-ACT maintains cooperative behavior under lower gain conditions ($r=4.0$). This stems from the Actor-Critic networks after Phase 1 training, which enhances spatial synergy effects and biases agents toward cooperative choices. Comparative cross-algorithm experiments reveal PPO-ACT’s superior critical behavioral characteristics: the minimum enhancement factor $r$ required to trigger cooperative emergence is significantly reduced, while the $r$ threshold for sustaining stable cooperation also decreases accordingly. Our PPO-ACT framework overcomes the inherent limitation of standard PPO where randomly initialized actor-critic networks fail to discover cooperative strategies under low enhancement factors ($r<5.1$). Through curriculum learning that implements phased reward shaping (Phase 1: $r=5.0$ Phase 2: $r=4.0$), the algorithm progressively guides policy networks to develop cooperation-enabling representations. This stands in sharp contrast to conventional PPO’s behavior at $r=4.0$, where the system inevitably converges to all-defector outcomes. In such cases, random initialization of policy parameters traps agents in defection-dominated Nash equilibria, completely preventing cooperation emergence. PPO-ACT and PPO exhibit typical bistable convergence patterns - systems exclusively reach all-cooperator or all-defector states, with extremely low probabilities of mixed cooperator-defector coexistence. This convergence behavior demonstrates nonlinear amplification effects inherent in the policy update mechanism, where minor initial strategy deviations become exponentially amplified through reinforcement learning, ultimately driving the system to phase transition into pure-strategy steady states. Q-learning exhibits a gradual improvement pattern in cooperation levels. However, even with high enhancement factors ($r=6.0$), the algorithm fails to achieve cooperation fractions exceeding $60\%$. This performance ceiling directly reflects the fundamental constraints of discrete Q-tables, which cannot effectively model spatial correlations between agents. Although the Fermi update rule can generate cooperative behavior when $r\geq 3.7$, it requires $r\geq 5.1$ to achieve stable full cooperation, demonstrating the inefficiency of local update rules. The experimental results show that PPO-ACT enables cooperators to persist under more stringent gain conditions. Our model offers crucial design principles for developing spatially adaptive multi-agent architectures.
3.7 Spatial strategy structures evolved by PPO-ACT
Fig. 7 shows the stable spatial strategy distribution formed by PPO after 1,000 training iterations under critical parameter conditions ($r=4.0$). The initialization strategy places defectors in the upper half and cooperators in the lower half. Immediate rewards (marked in red) reveal fundamental dynamical mechanisms in spatial games.
Figure 7: 20$\times$20 grids were cropped from PPO-ACT’s strategy matrix snapshots exhibiting cooperator (white) and defector (black) distributions. Red numeric annotations denote agent payoff values. The initialization strategy places defectors in the upper half and cooperators in the lower half. $r=4.0$.
Cooperator clusters exhibit characteristic defensive structures. Central cooperators achieve maximum rewards ($+18.0$), verifying intra-cluster synergistic effects. Edge cooperators lose $2.4$ per adjacent defector. This reward gradient precisely reflects spatial positioning’s critical impact on strategy effectiveness. Defectors show clear contact-dependent payoff patterns and adjacent to cooperators achieve maximum exploitation gains ($+19.2$). This spatial dependency demonstrates that the effectiveness of defection strategies is wholly contingent upon neighboring contact opportunities. Through policy gradient optimization, PPO-ACT spontaneously forms stable cooperative-defective phase separation. Cooperators minimize edge exposure through tight clustering, while defectors dynamically adjust positions to maintain exploitation opportunities. The spatial self-organization process produces distinct behavioral patterns. Within cooperative clusters, increased reward levels effectively reinforce collective defensive mechanisms. Meanwhile, restricted contact opportunities for defectors drive the spontaneous emergence of exploitation frontiers surrounding cooperator groups. This phenomenon provides algorithmic-level micro-mechanisms for network reciprocity theory. PPO-ACT demonstrates significant advantages at low enhancement levels ($r=4.0$). The algorithm empowers agents to acquire spatial optimization capabilities through policy gradient methods independently. This innovative approach effectively addresses the constraints inherent in conventional imitation-based dynamics. This learning capacity provides new empirical evidence for studying cooperation evolution in multi-agent systems. Notably, PPO-ACT agents not only learn cooperative strategies but master higher-order skills of enhancing effectiveness through spatial configuration.
@@ 3.8 Hyperparameter sensitivity analysis of PPO-ACT
Figure 8: Impact of initial learning rate $\alpha$ on results.
The discount factor $\gamma$ is a crucial hyperparameter in PPO-ACT that determines the agent’s consideration of future rewards. In this experiment, $\rho=0.01$. As shown in Fig. 9, experimental results demonstrate that at $\gamma=0.99$ (PPO’s default), PPO-ACT exhibits strongest cooperation promotion, achieving stable cooperation at minimal $r$ values. This suggests higher $\gamma$ values better evaluate long-term cooperative benefits in collaborative tasks.
Figure 9: Impact of discount factor $\gamma$ on results.
@@
Figure 10: Impact of discount factor $\delta$ on results.
The entropy weight $\rho=0.001$ ensures stable policy transfer by balancing exploration and exploitation. Higher $\rho$ values hinder convergence to cooperation as excessive exploration disrupts learned strategies during Phase 2. This carefully tuned low entropy promotes reliable cooperation at challenging reward levels while maintaining necessary adaptability.
Figure 11: Impact of discount factor $\rho$ on results.
@@ 4 Conclusions
Despite these achievements, several limitations remain. The computational complexity and sensitivity to hyperparameter settings require further improvement. Additionally, scalability in large-scale systems needs verification. Future studies will extend PPO-ACT applications to more complex network structures. Additional work should examine cooperative behaviors in populations with heterogeneous agents. These efforts will expand both the theoretical depth and practical applications of this work.

### v2

@@ 3.2 PPO-ACT with half-and-half initialization
Figure 2: Evolution of cooperative behavior in PPO-ACT under spatial heterogeneity. (a,c) Temporal evolution under enhancement factors $r=3.0$ (a) and $r=4.0$ (c) showing the complete Phase 1 + Phase 2 training process, with strategy space snapshots (white: cooperator, black: defector) at iterations $\{0,10,100,1000,10000\}$. (b,d) Corresponding Phase 2-only results for $r=3.0$ (b) and $r=4.0$ (d) with snapshots at $\{0,1,10,100,1000\}$ iterations. Initial conditions: defectors (red) occupy the upper grid half, cooperators (blue) lower half. All panels share identical color mapping and spatial scale bars. The initial training phase with r=5.0 established a cooperative foundation, enabling agents to partially overcome payoff disadvantages in the second phase. Evolutionary trajectories in panels (c) and (d) demonstrate that at r=3.0, the critically low cooperation payoff leads to complete dominance by defectors, while at r=4.0, the population rapidly converges to universal cooperation.
@@ 3.3 PPO-ACT with bernoulli random initialization
(b) Phase 2
Figure 3: Evolution of cooperative behavior under Bernoulli initialization ($p=0.5$). (a) Complete training process with Phase 1 ($r=5.0$, 0-999 iterations) and Phase 2 ($r=4.0$, from iteration 1000). Snapshots captured at $t=[0,10,100,1000,10000]$ (white: cooperator, black: defector). (b) Phase 2-only dynamics after strategy reinitialization. Snapshots at $t=[0,1,10,100,1000]$ demonstrate immediate cooperation fixation. Bernoulli initialization ($p=0.5$) enables faster cooperative convergence than spatially segregated setups, with Phase 2 achieving instant ($t=1$) global cooperation fixation.
@@ 3.4 PPO-ACT with all-defectors initialization
Figure 4: Initial agent strategies were set with all agents as defectors. (a) Full process: Phase 1 ($r=5.0$, 0-999 iterations) and Phase 2 ($r=4.8$, from iteration 1000). Snapshots at $t=[0,10,100,1000,10000]$. (b) Phase 2-only process. Snapshots at $t=[0,1,10,100,1000]$. White indicates cooperators, and black denotes defectors. PPO-ACT can achieve global cooperation from all-defector initializations, though requiring higher enhancement factors $r$ compared to stochastic or spatially mixed initial conditions to reach stable cooperative equilibria.
@@ 3.5 Comparative analysis of algorithms
Figure 5: Comparative analysis of PPO-ACT, PPO, Q-learning, and Fermi update rule: Temporal evolution curves of cooperators (blue) and defectors (red) at $r=4.0$, with corresponding snapshots of cooperators (white) and defectors (black). The initialization strategy places defectors in the upper half and cooperators in the lower half. Snapshots are shown chronologically from left to right (0, 10, 100, 1000, and 10000 iterations). Compared to other algorithms, our PPO-ACT can converge to a state where all agents are cooperators at a lower enhancement factor $r$.
@@
3.6 Statistical analysis of PPO-ACT
To verify the reliability of PPO-ACT and quantitatively evaluate the enhancement from the ACT module, we systematically tested enhancement factors $r\in[3.0,6.0]$ in 50 independent trials per value. Results are presented through three complementary statistical visualizations.
Error Bar Analysis (Fig. 6) demonstrates the fundamental distinction between PPO-ACT and baseline PPO in terms of mean cooperation rates and their variability. Both algorithms exhibit binary outcomes (either 0% or 100% cooperation) in individual runs, resulting in substantial standard deviations across all $r$ values. Specifically, PPO-ACT achieves cooperation at lower $r$ values compared to PPO. It confirms the ACT’s effectiveness in lowering the cooperation barrier.
(a) PPO-ACT
(b) PPO
Figure 6: Comparison of mean cooperation rates and standard deviation between PPO-ACT and PPO. Due to the effect of ACT, our PPO-ACT exhibits cooperative behavior at a smaller $r$ than PPO.
Violin plot analysis (Fig. 7) reveals key findings through median, mean, and distribution examination. Compared with PPO at $r=5.0$, when the enhancement factor $r\in[3.0,5.0]$, PPO-ACT demonstrates significantly superior performance in both median cooperation rates and mean values. Moreover, Fig. 7) visualizes the probability density distribution of cooperation rates across different r values. It clearly shows that PPO-ACT achieves earlier emergence and more stable high-density cooperative regions in its distribution.
(a) PPO-ACT
(b) PPO
Figure 7: Probability density of cooperation fractions for PPO-ACT (left) and PPO (right) across enhancement factors $r\in[3.0,6.0]$. The violin plot indicates that PPO-ACT has a higher probability of converging to a state where all agents are collaborators even when r is small.
As shown in Table 2, PPO-ACT and PPO exhibit distinct behavioral patterns in the spatial public goods game. PPO-ACT demonstrates progressive evolutionary characteristics, with its cooperation rate confidence intervals evolving smoothly from 0.48–0.76 at $r=3.0$ to 0.84–1.00 at $r=4.7$, reflecting the probabilistic decision-making nature of this strategy. In contrast, PPO displays a clear threshold transition behavior: maintaining complete defection when $r<5.0$, transitioning to partial cooperation at $r=5.0$, and achieving full cooperation when $r>5.0$. The confidence intervals marked as ‘nan’ in the table indicate conditions where the strategy outputs were completely deterministic (e.g., PPO-ACT’s full cooperation at $r\geq 4.9$), making valid confidence interval computation impossible. The 0.00–0.00 intervals represent cases where the observed cooperation rates were extremely low but not absolutely zero. These results suggest that PPO-ACT is better suited for complex social dilemma scenarios, while PPO’s deterministic strategy shows significant limitations in flexibility.
Table 2: $95\%$ confidence intervals comparison for cooperation fractions
r 3.0 3.1 3.2 3.3 3.4 3.5 3.6
PPO-ACT $0.48$ – $0.76$ $0.34$ – $0.62$ $0.26$ – $0.54$ $0.34$ – $0.62$ $0.64$ – $0.88$ $0.52$ – $0.80$ $0.42$ – $0.70$
PPO $0.00$ – $0.00$ nan – nan nan – nan nan – nan $0.00$ – $0.00$ nan – nan $0.00$ – $0.00$
r 3.7 3.8 3.9 4.0 4.1 4.2 4.3
PPO-ACT $0.44$ – $0.72$ $0.44$ – $0.72$ $0.42$ – $0.70$ $0.66$ – $0.90$ $0.50$ – $0.78$ $0.66$ – $0.90$ $0.55$ – $0.81$
PPO nan – nan nan – nan nan – nan $0.00$ – $0.00$ nan – nan nan – nan $0.00$ – $0.00$
r 4.4 4.5 4.6 4.7 4.8 4.9 5.0
PPO-ACT $0.71$ – $0.93$ $0.59$ – $0.85$ $0.79$ – $0.97$ $0.84$ – $1.00$ $0.76$ – $0.96$ nan – nan nan – nan
PPO $0.00$ – $0.00$ nan – nan nan – nan nan – nan $0.00$ – $0.00$ $0.00$ – $0.00$ $0.26$ – $0.54$
@@ 3.7 Comparative analysis of algorithms across different $r$
PPO-ACT maintains cooperative behavior under lower gain conditions ($r=4.0$). This stems from the Actor-Critic networks after Phase 1 training, which enhances spatial synergy effects and biases agents toward cooperative choices. Comparative cross-algorithm experiments reveal PPO-ACT’s superior critical behavioral characteristics: the minimum enhancement factor $r$ required to trigger cooperative emergence is significantly reduced, while the $r$ threshold for sustaining stable cooperation also decreases accordingly. Our PPO-ACT framework overcomes the inherent limitation of standard PPO where randomly initialized actor-critic networks fail to discover cooperative strategies under low enhancement factors ($r<5.1$). Through curriculum learning that implements phased reward shaping (Phase 1: $r=5.0$ Phase 2: $r=4.0$), the algorithm progressively guides policy networks to develop cooperation-enabling representations. This stands in sharp contrast to conventional PPO’s behavior at $r=4.0$, where the system inevitably converges to all-defector outcomes. In such cases, random initialization of policy parameters traps agents in defection-dominated Nash equilibria, completely preventing cooperation emergence.
@@
Figure 8: Performance comparison of PPO-ACT, PPO, Q-learning, and Fermi update rule across different r-values: cooperators (blue) and defectors (red). Initial agent strategies were set with all defectors in the upper grid half and all cooperators in the lower half. Compared to other algorithms, starting from a lower enhancement factor $r$, our PPO-ACT can stably converge to a state where all agents’ strategies are cooperative strategies.
@@
PPO-ACT and PPO exhibit typical bistable convergence patterns - systems exclusively reach all-cooperator or all-defector states, with extremely low probabilities of mixed cooperator-defector coexistence. This convergence behavior demonstrates nonlinear amplification effects inherent in the policy update mechanism, where minor initial strategy deviations become exponentially amplified through reinforcement learning, ultimately driving the system to phase transition into pure-strategy steady states. Q-learning exhibits a gradual improvement pattern in cooperation levels. However, even with high enhancement factors ($r=6.0$), the algorithm fails to achieve cooperation fractions exceeding $60\%$. This performance ceiling directly reflects the fundamental constraints of discrete Q-tables, which cannot effectively model spatial correlations between agents. Although the Fermi update rule can generate cooperative behavior when $r\geq 3.7$, it requires $r\geq 5.1$ to achieve stable full cooperation, demonstrating the inefficiency of local update rules. The experimental results show that PPO-ACT enables cooperators to persist under more stringent gain conditions. Our model offers crucial design principles for developing spatially adaptive multi-agent architectures.
@@ 3.8 Hyperparameter sensitivity analysis of PPO-ACT
Figure 9: Impact of initial learning rate $\alpha$ on results. In this experiment, the model performed the best when $\alpha=0.001$.
The discount factor $\gamma$ is a crucial hyperparameter in PPO-ACT that determines the agent’s consideration of future rewards. In this experiment, $\rho=0.01$. As shown in Fig. 10, experimental results demonstrate that at $\gamma=0.96$, PPO-ACT exhibits strongest cooperation promotions. Therefore, the $\gamma$ of Phase 2 is set to 0.96.
Figure 10: Impact of discount factor $\gamma$ on results. Experimental data indicate $\gamma=0.96$ as the parameter value yielding maximum model efficacy.
@@
Figure 11: Impact of discount factor $\delta$ on results. The model reaches its performance peak when $\delta$ is set to 0.96.
The entropy weight $\rho=0.001$ ensures stable policy transfer by balancing exploration and exploitation (see Fig. 12). Higher $\rho$ values hinder convergence to cooperation as excessive exploration disrupts learned strategies during Phase 2. This carefully tuned low entropy promotes reliable cooperation at challenging reward levels while maintaining necessary adaptability.
Figure 12: Impact of discount factor $\rho$ on results. The optimal model performance was achieved at $\rho=0.001$
@@ 4 Conclusions
Despite these achievements, several limitations remain. The computational complexity and sensitivity to hyperparameter settings require further improvement. Additionally, scalability in large-scale systems needs verification. Future studies will extend PPO-ACT applications to more complex network structures. Additional work should examine cooperative behaviors in populations with heterogeneous agents. Furthermore, incorporating language-based utility functions [58] could enhance the model’s ability to simulate human-AI interaction scenarios, where linguistic framing plays a critical role in decision-making. These efforts will expand both the theoretical depth and practical applications of this work.